Ticket Number: HEA-NOC/20051026-2 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20051026 10:24 UTC+1 Problem Start: 20051026 10:00 UTC+1
Ticket Closed: 20051118 12:15 UTC Problem End: 20051118 12:10 UTC
Site/line: HEA-NOC, BH-Citywest GE (HEAN0015), BH-Citywest GE
(HEAN0104), Bluenet BH-Citywest
The 2 BT Gig/E links from Brooklawn House to Citywest will be ceased at the
end of this month. To replace these, 2 Gig/E circuits will be provisioned
over the existing 10Gig dark fibre connection from Brooklawn House to
No customer impact is expected. Both sets of circuits will co-exist before
the BT circuits cease.
20051026 09:57 OB Cabling in Brooklawn house complete; scheduling patching
in Citywest for tomorrow. Configs have been specified and provisioning of
layer 2 circuits between Brooklawn House & Citywest will occur this
afternoon. Layer 3 configs will be applied (with suitable OSPF recosting
applied during "dual period") tomorrow afternoon.
20051028 16:33 OB One provisioned circuit from Brooklawn House to Citywest
is complete. Suspected SFP module in Citywest is delaying provisioning of
2nd circuit. This module will be replaced next week. In the meantime, BT
circuits remain in place.
20051101 16:45 DC Suspect SFP replacement will take place tomorrow. One of
the BT links will be deprioritised on Thursday, 3rd of November at 9:00,
causing traffic to flow via the working Dark Fibre link.
20051102 13:53 EK Donal managed to isolate a fault fibre patch from
Hyperion to the patch panel. The patch has been replaced and the link is
20051107 14:51 OB Provisioning of 2nd circuit is scheduled for tomorrow
morning at 9am. Upon completion, costs on the 2nd BT link will be
increased causing traffic to flow over this newly provisioned circuit.
20051108 12:27 OB Costs increased on BT circuit and traffic is now flowing
as expected over 2 dark fibre circuits. (This circuit was actually
provisioned last week) Will continue to closely monitor for 24 hours.
20051118 12:14 OB Links have been stable for over 1 week. Closing ticket.
This ticket can be monitored at http://www.hea.net/tickets/20051026-2